# Understanding Oversmoothing in Diffusion-Based GNNs From the Perspective of Operator Semigroup Theory

### v1

@@ 7 Conclusions
Fix $h>0$ and consider the h-skeleton $\mathbf{Z}_{n}:=\mathbf{X}_{nh}$ of $\mathbf{X}$. Since
$\mathbb{P}(\mathbf{Z}_{n+1}=u_{n+1}|\mathbf{Z}_{0}=u_{0},\ldots,\mathbf{Z}_{n}=u_{n})=p(h,u_{n},u_{n+1}),$
h-skeleton chain $\mathbf{Z}_{n}$ is discrete time Markov chain with transition matrix $P(h)$. Since $\mathcal{G}$ is a finite connected graph, for all $u,v\in\mathcal{V}$, there exists some $t$, such that $p(t,u,v)>0$. Therefore $\mathbf{Z}_{n}$ is irreducible and aperiodic. From classic result of discrete time Markov chain(Norris, 1998), there exists a unique invariant measure $\mu$, for all $u,v\in\mathcal{V}$
@@
and find $N\in\mathbb{N}$ such that
$|p(nh,u,v)-\mu(v)|<\frac{\epsilon}{2},\quad n>N.$
For $t>Nh$ we have $nh\leq t<(n+1)h$ for some $n>N$ and
@@
Hence
$p(t,u,v)\rightarrow\mu(v),\quad t\rightarrow\infty.$
B.3 Proof of Theorem 5.3
@@
$=\mathbb{P}(\mathbf{X}_{t+s}=u,\tau>t+s\;|\;\mathbf{X}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}$
$=\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=u\;|\;\tilde{\mathbf{X}}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}.$
@@
$=\mathbf{1}_{\tilde{\mathbf{X}}_{s}=\partial}+\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=\partial\;|\;\tilde{\mathbf{X}}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}$
$=\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=\partial\;|\;\tilde{\mathbf{X}}_{s}).$
@@
Lemma B.1 (The Feynman-Kac formula, Theorem 8.2.1 of Oksendal (2013)).
Let $f\in C_{0}^{2}(\mathbb{R}^{n})$ and $q\in C(\mathbb{R}^{n})$. Assume that $q$ is lower bounded. Put
$v(t,x)=\mathbb{E}^{x}\left[\exp\left(-\int_{0}^{t}q(\xi_{s})\mathrm{d}s\right)f(\xi_{t})\right].$
Then
$\frac{\partial v}{\partial t}=Av-qv,\quad v(0,x)=f(x),\quad t>0,\;x\in\mathbb{R}^{n},$
where $A$ is generator of diffusion process $\xi$.
@@
where $c(u)=-\lim\limits_{t\rightarrow 0^{+}}\frac{1}{t}\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$. Then, by lemma B.1,
@@
That is, the generator of $\tilde{\mathbf{X}}$ is
$\tilde{\mathcal{A}}=\mathcal{A}+c(u).$
Appendix C Experimental Details and Further Results
C.1 Configurations and Results for Oversmoothing
Synthetic Cora Dataset The Synthetic Cora dataset, as introduced by Zhu et al. (2020), is a variant of the well-known Cora dataset tailored for graph neural network (GNN) experiments. This dataset is characterized by its homophily index, which ranges from 0 to 1. Homophily in this context refers to the tendency of nodes in the graph to be connected to other nodes with similar features or labels. A homophily index of 0 indicates no homophily (i.e., connections are completely independent of node labels), while an index of 1 signifies perfect homophily (i.e., nodes are only connected to others with the same label). This range allows for controlled experiments to understand the behavior of GNNs under varying degrees of homophily, which is crucial for assessing the robustness and adaptability of different models. We provide the detailed information of the synthetic Cora networks in Table 2.
Table 2: Detailed characteristics of the Synthetic Cora dataset across various homophily levels. Each homophily level maintains identical counts of features (1,433), nodes (1,480), edges (5,936), and classes (5), ensuring consistent structural properties while varying node similarity.
Homophily Avg. Degree Max. Degree Min. Degree
0.0 3.98 84.33 1.67
0.1 3.98 71.33 2.00
0.2 3.98 73.33 1.67
0.3 3.98 70.00 2.00
0.4 3.98 77.67 2.00
0.5 3.98 76.33 2.00
0.6 3.98 76.00 1.67
0.7 3.98 67.67 2.00
0.8 3.98 58.00 1.67
0.9 3.98 58.00 1.67
1.0 3.98 51.00 2.00
Configurations For Figure 1 and Further Results For all ergodicity-breaking terms, we use a first-order truncation for ergodicity-breaking terms within our proposed models: Ours-Exp, Ours-ExpN, Ours-Log, Ours-Sin, and Ours-Cos. We further show the comprehensive results of oversmoothing with different homophily levels and ergodicity-breaking orders in Figure 2.
(a) Dirichlet Energy Demonstration For Attention Block With Training
(b) Dirichlet Energy Demonstration For Constant Block With Training
(c) Dirichlet Energy Demonstration For Attention Block Without Training
(d) Dirichlet Energy Demonstration For Constant Block Without Training
Figure 2: Comprehensive Demonstration of the Dirichlet energy for extensive diffusion-based GNNs, including GRAND-l, GRAND-nl, and our method (Exp, Expn, Log, Sin, Cos) with different orders {1,2,3,4} over increasing network depths and homophily levels on synthetic Cora dataset.
C.2 Configurations for Node Classification Tasks
@@
The ergodicity-breaking terms introduced in Sec. 6.3 were configured with specific truncated orders N. The optimal hyperparameters for each term were meticulously tuned based on the performance of the GREAD (Choi et al., 2023). The optimal configurations are systematically listed in Tables 3-7. Each table corresponds to one of the ergodicity-breaking terms and includes the tuned parameters.
Table 3: Best hyperparameters of Exponential Ergodicity Breaking Terms
Hyperparameters Texas Wisconsin Cornell Film Squirrel Chameleon Cora Citeseer PubMed
Order 1 1 2 2 1 3 2 1 2
Block Attention Constant Constant Constant Constant 1 1 2 3
$\alpha$ SC SC SC VC VC VC SC VC SC
$\beta$ SC SC SC VC VC VC SC SC VC
learning rate 0.0100 0.0154 0.0082 0.0079 0.0171 0.0068 0.0105 0.0024 0.0108
weight decay 0.0247 0.0090 0.0280 0.0014 0.0000 0.0060 0.0146 0.0005
input dropout 0.47 0.54 0.49 0.42 0.52 0.68 0.53 0.50 0.36
dropout 0.48 0.48 0.32 0.65 0.09 0.05 0.45 0.47 0.26
dim(H) 256 256 64 64 128 256 128 256 128
step size $\tau$ 1.0 0.24 0.25 0.1 0.72 1.72 0.5 0.45 0.1
time $T$ 1.461 1.85 0.115 0.2 5.725 1.91 3.575 2.28 1
ODE solver Euler Euler RK4 Euler Euler Euler RK4 Euler RK4
Table 4: Best hyperparameters of Negative exponential Ergodicity Breaking Term
Hyperparameters Texas Wisconsin Cornell Film Squirrel Chameleon Cora Citeseer PubMed
Order 1 1 2 3 1 3 1 3 1
Block Attention Attention Attention Attention Constant 1 1 2 3
$\alpha$ SC SC VC VC VC VC SC VC SC
$\beta$ SC VC VC SC VC VC SC VC VC
learning rate 0.0100 0.0154 0.0082 0.0079 0.0171 0.0068 0.0105 0.0024 0.0108
weight decay 0.0247 0.0090 0.0280 0.0014 0.0000 0.0060 0.0146 0.0005
input dropout 0.47 0.54 0.49 0.42 0.52 0.68 0.53 0.50 0.36
dropout 0.48 0.48 0.32 0.65 0.09 0.05 0.45 0.47 0.26
dim(H) 64 256 64 32 256 256 128 128 64
step size $\tau$ 1.0 0.2 0.2 0.1 0.94 1.66 0.25 0.5 0.8
time $T$ 1.461 1.6 0.135 0.35 5.9 1.9 3.49 2.38 1.74
ODE solver Euler RK4 RK4 Euler Euler RK4 RK4 Euler RK4
Table 5: Best hyperparameters of Logarithm Ergodicity Breaking Term
Hyperparameters Texas Wisconsin Cornell Film Squirrel Chameleon Cora Citeseer PubMed
Order 2 2 3 1 1 1 2 2 2
Block Constant Constant Constant Constant Constant 1 1 2 3
$\alpha$ SC SC SC VC VC VC VC VC SC
$\beta$ SC SC VC VC VC VC SC VC SC
learning rate 0.0100 0.0154 0.0082 0.0079 0.0171 0.0068 0.0105 0.0024 0.0108
weight decay 0.0247 0.0090 0.0280 0.0014 0.0000 0.0060 0.0146 0.0005
input dropout 0.47 0.54 0.49 0.42 0.52 0.68 0.53 0.50 0.36
dropout 0.48 0.48 0.32 0.65 0.09 0.05 0.45 0.47 0.26
dim(H) 64 128 256 32 128 256 64 64 128
step size $\tau$ 1.0 0.22 0.24 0.16 0.8 1.62 0.5 0.45 0.8
time $T$ 1.461 1.8 0.13 0.45 5.925 1.99 3.79 2.26 1.74
ODE solver Euler RK4 Euler RK4 Euler Euler RK4 Euler RK4
Table 6: Best hyperparameters of Sine Ergodicity Breaking Term
Hyperparameters Texas Wisconsin Cornell Film Squirrel Chameleon Cora Citeseer PubMed
Order 3 3 3 1 1 1 1 2 3
Block Constant Attention Constant Attention Constant 1 1 2 3
$\alpha$ SC SC VC SC SC VC VC VC SC
$\beta$ VC VC VC VC VC VC SC VC VC
learning rate 0.0100 0.0154 0.0082 0.0079 0.0171 0.0068 0.0105 0.0024 0.0108
weight decay 0.0247 0.0090 0.0280 0.0014 0.0000 0.0060 0.0146 0.0005
input dropout 0.47 0.54 0.49 0.42 0.52 0.68 0.53 0.50 0.36
dropout 0.48 0.48 0.32 0.65 0.09 0.05 0.45 0.47 0.26
dim(H) 32 64 32 128 128 256 128 256 256
step size $\tau$ 1.0 0.22 0.21 0.16 0.98 1.66 0.5 0.45 0.8
time $T$ 1.461 1.875 0.125 0.3 5.925 1.99 3.57 2.24 1.74
ODE solver Euler Euler Euler RK4 Euler Euler RK4 Euler RK4
Table 7: Best hyperparameters of Cosine Ergodicity Breaking Term
Hyperparameters Texas Wisconsin Cornell Film Squirrel Chameleon Cora Citeseer PubMed
Order 1 1 2 2 1 3 1 2 3
Block Attention Attention Attention Attention Constant 1 1 2 3
$\alpha$ SC SC VC VC SC SC VC VC VC
$\beta$ VC SC VC SC VC VC VC VC SC
learning rate 0.0100 0.0154 0.0082 0.0079 0.0171 0.0068 0.0105 0.0024 0.0108
weight decay 0.0247 0.0090 0.0280 0.0014 0.0000 0.0060 0.0146 0.0005
input dropout 0.47 0.54 0.49 0.42 0.52 0.68 0.53 0.50 0.36
dropout 0.48 0.48 0.32 0.65 0.09 0.05 0.45 0.47 0.26
dim(H) 64 64 32 64 128 256 128 256 64
step size $\tau$ 1.0 0.26 0.16 0.1 0.98 1.74 0.25 0.5 0.8
time $T$ 1.44 1.625 0.12 0.125 5.625 1.71 3.98 2.24 1.74
ODE solver Euler RK4 RK4 RK4 Euler RK4 Euler RK4 RK4

### v2

@@ B.2. Proof of Theorem 2
Fix $h>0$ and consider the h-skeleton $\mathbf{Z}_{n}:=\mathbf{X}_{nh}$ of $\mathbf{X}$. Since $\mathbb{P}(\mathbf{Z}_{n+1}=u_{n+1}|\mathbf{Z}_{0}=u_{0},\ldots,\mathbf{Z}_{n}=u_{n})=p(h,u_{n},u_{n+1}),$ h-skeleton chain $\mathbf{Z}_{n}$ is discrete time Markov chain with transition matrix $P(h)$. Since $\mathcal{G}$ is a finite connected graph, for all $u,v\in\mathcal{V}$, there exists some $t$, such that $p(t,u,v)>0$. Therefore $\mathbf{Z}_{n}$ is irreducible and aperiodic. From classic result of discrete time Markov chain(Norris, 1998), there exists a unique invariant measure $\mu$, for all $u,v\in\mathcal{V}$
@@
and find $N\in\mathbb{N}$ such that $|p(nh,u,v)-\mu(v)|<\frac{\epsilon}{2},n>N.$ For $t>Nh$ we have $nh\leq t<(n+1)h$ for some $n>N$ and
@@
Hence $p(t,u,v)\rightarrow\mu(v),\quad t\rightarrow\infty.$
B.3. Proof of Theorem 3
@@
$=\mathbb{P}(\mathbf{X}_{t+s}=u,\tau>t+s\;|\;\mathbf{X}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}=\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=u\;|\;\tilde{\mathbf{X}}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}.$
@@
$=\mathbf{1}_{\tilde{\mathbf{X}}_{s}=\partial}+\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=\partial\;|\;\tilde{\mathbf{X}}_{s})\mathbf{1}_{\tilde{\mathbf{X}}_{s}\neq\partial}=\mathbb{P}(\tilde{\mathbf{X}}_{t+s}=\partial\;|\;\tilde{\mathbf{X}}_{s}).$
@@
Lemma 1 (The Feynman-Kac formula, Theorem 8.2.1 of Oksendal (2013)).
Let $f\in C_{0}^{2}(\mathbb{R}^{n})$ and $q\in C(\mathbb{R}^{n})$. Assume that $q$ is lower bounded. Put $v(t,x)=\mathbb{E}^{x}\left[\exp\left(-\int_{0}^{t}q(\xi_{s})\mathrm{d}s\right)f(\xi_{t})\right].$ Then $\frac{\partial v}{\partial t}=Av-qv,\quad v(0,x)=f(x),\quad t>0,\;x\in\mathbb{R}^{n},$ where $A$ is generator of diffusion process $\xi$.
@@
where $c(u)=-\lim\limits_{t\rightarrow 0^{+}}\frac{1}{t}\mathbb{P}(\tau\leq t|\mathbf{X}_{0}=u)$. Then, by lemma 1,
@@
That is, the generator of $\tilde{\mathbf{X}}$ is $\tilde{\mathcal{A}}=\mathcal{A}+c(u).$
Appendix C Experimental Details
C.1. Configurations for Oversmoothing
Synthetic Cora Dataset. The Synthetic Cora dataset, as introduced by Zhu et al. (2020), is a variant of the well-known Cora dataset tailored for graph neural network (GNN) experiments. This dataset is characterized by its homophily index, which ranges from 0 to 1.
Configurations For Figure 1. For all ergodicity-breaking terms, we use a first-order truncation for ergodicity-breaking terms within our proposed models: Ours-Exp, Ours-ExpN and Ours-Log.
C.2. Configurations for Node Classification Tasks
@@
The optimal hyperparameters for each term were meticulously tuned based on the settings in GREAD (Choi et al., 2023). The ergodicity-breaking terms introduced in Sec. 7.3 were configured with specific truncated orders N. Empirically, we observe that the hyperparameters specified in GREAD demonstrate robust experimental performance. The code repository for the paper is available at: https://github.com/LOGO-CUHKSZ/SGOS